# Quantum Learning and Estimation for Distribution Networks and Energy Communities Coordination

### v1

@@ V Conclusion
In this paper, a quantum learning and estimation approach is proposed to address the bilevel DNs-ECs coordination problem. By leveraging quantum properties including quantum superposition and entanglement, the proposed Q-TCN-LSTM model can efficiently learn the incentive-response characteristics of ECs, realizing accurate estimation of user electricity consumption behavior under limited information. Additionally, QAE is introduced to facilitate evaluations with numerous scenarios, achieving a quadratic speedup over classical MC methods, significantly reducing computational resource consumption and execution time. In this way, the proposed quantum learning and estimation approach efficiently facilitates the coordination between DNs and ECs to generate better pricing strategies to regulate the net load profiles of ECs, thereby reducing the operational risk introduced to the DNs. As illustrated by the case studies, the proposed quantum-enhanced approach demonstrates superior performance in terms of both mapping accuracy and computational efficiency.
Future work will consider integrating quantum generative adversarial networks for scenario generation into the quantum estimation process to further enhance the efficiency of estimation and optimization processes.
-A Operation Model of Multi-Type ECs
Different types of ECs exhibit heterogeneous characteristics in terms of device composition and operational preferences, leading to diverse operational behaviors.
-A1 Residential ECs
Residential ECs (e.g., individual households, apartment complexes, and community clusters) typically integrate rooftop PV systems, electric vehicles (EVs), inflexible loads (e.g., essential household appliances), and flexible loads (e.g., air conditioners). Flexible loads are modeled with both temporal shiftability and power reducibility. EVs function as energy storage (ES) to provide ancillary service through smart charging/discharging. However, due to daily commuting patterns, EVs are typically unavailable during daytime hours and can only participate in energy management during night. Additionally, charging and discharging cycles are constrained to mitigate battery degradation.
The objective of residential ECs operation is to minimize the total operational cost. Thus, the overall operation model of residential ECs is formulated as follows:
$\min$ $\sum_{t=1}^{T}\pi^{\text{EC}}_{t}P^{\text{EX,R}}_{t}\Delta t,$ (16a)
s.t. $0\leq P^{\text{E,c}}_{t}\leq D^{\text{E,c}}_{t}\overline{P}^{\text{E}},\ 0\leq P^{\text{E,d}}_{t}\leq D^{\text{E,d}}_{t}\overline{P}^{\text{E}},\ \forall t\in\mathcal{T},$ (16b)
$0\leq D^{\text{E,c}}_{t}+D^{\text{E,d}}_{t}\leq 1,\ \forall t\in\mathcal{T},$ (16c)
$SoC_{t+1}=SoC_{t}+\Delta t(P_{t}^{\mathrm{E},\mathrm{c}}\eta^{\mathrm{c}}-P_{t}^{\mathrm{E},\mathrm{d}}/\eta^{\mathrm{d}})/E,\ \forall t\in\mathcal{T},$ (16d)
$\underline{SoC}\leq SoC_{t}\leq\overline{SoC},\ \forall t\in\mathcal{T},$ (16e)
$SoC_{t_{1}}=\overline{SoC},\ SoC_{t_{2}}=\underline{SoC},$ (16f)
$P^{\text{E,c}}_{t}=0,\ P^{\text{E,d}}_{t}=0,\ \forall t\notin\mathcal{T},$ (16g)
$0\leq\sum_{t\in\mathcal{T}}(D^{\text{E,c}}_{t}+D^{\text{E,d}}_{t})\leq N^{\text{E}},$ (16h)
$-\rho\hat{P}^{\text{L}}_{t}\leq P^{\text{L,s}}_{t}\leq\rho\hat{P}^{\text{L}}_{t},$ (16i)
$0\leq P_{t}^{\text{R,c}}\leq\hat{P}_{t}^{\text{R}},$ (16j)
$P^{\text{EX,R}}_{t}\!=\!-(\hat{P}_{t}^{\text{R}}-P_{t}^{\text{R,c}})\!+\!\hat{P}^{\text{L}}_{t}\!+\!P^{\text{L,s}}_{t}\!+\!P^{\text{E,c}}_{t}\!-\!P^{\text{E,d}}_{t},$ (16k)
$\underline{P}^{\text{EX,R}}_{t}\leq P^{\text{EX,R}}_{t}\leq\overline{P}^{\text{EX,R}}_{t},$ (16l)
where (16b)-(16c) constrain the charging/discharging power (i.e., $P^{\text{E,c}}_{t}/P^{\text{E,d}}_{t}$) limited by $\overline{P}^{\text{E}}$ and state (i.e., $D^{\text{E,c}}_{t}/D^{\text{E,d}}_{t}$) of ES. (16d)-(16f) further regulate the state of charge (SoC) dynamics with charging/discharging efficiency $\eta^{\text{c}}/\eta^{\text{d}}$ and ES capacity $E$. (16f) enforces that the SoC reaches its upper bound (i.e., $\overline{SoC}$) at the end of the nighttime (i.e., $t_{1}$), and its lower bound (i.e., $\underline{SoC}$) at the end of the daytime (i.e., $t_{2}$), and (16g) constrains the time periods (i.e., $\mathcal{T}=[1,t_{1}]\cup[t_{2},T]$) when ES is not allowed to charge/discharge. (16h) limits the total number of charging/discharging operations (i.e., $N^{\text{E}}$) to reduce battery degradation. (16i) constrains the flexible load shedding (i.e., $P^{\text{L,s}}_{t}$) of the forecasted load demand (i.e., $\hat{P}^{\text{L}}_{t}$). (16j) constrains the maximum energy curtailment (i.e., $P_{t}^{\text{R,c}}$) of the forecasted renewable energy generation (i.e., $\hat{P}_{t}^{\text{R}}$). (16k) is the power balance equation. $P^{\text{EX,R}}_{t}$ is the trading power of the residential EC limited by the upper bound $\overline{P}^{\text{EX}}_{t}$ in (16l). $\Delta t/T$ are the scheduling time interval/total time period. $\rho$ is the maximum shedding ratio of the flexible load reflecting the response willingness of the user.
-A2 Commercial ECs
Commercial ECs (e.g., shopping malls, office buildings, and hotels) are considered to be equipped with PV, ES, inflexible loads and flexible loads. Due to their customer-oriented operational nature, flexible loads in commercial ECs are typically restricted to curtailment rather than shifting. Such adjustments are often associated with customer discomfort or service quality degradation, which can lead to additional economic costs.
The objective of commercial ECs operation is to minimize the operational cost including the power trading cost and the flexible load adjustment cost. Thus, the overall operation model of commercial ECs is given as follows:
$\min$ $\sum_{t=1}^{T}\Delta t(\pi^{\text{EC}}_{t}P^{\text{EX,C}}_{t}+\alpha^{\text{L,s}}P^{\text{L,s}}_{t}),$ (17a)
[…]
$\sum_{t=1}^{T}(P_{t}^{\mathrm{E},\mathrm{c}}\eta^{\mathrm{c}}-P_{t}^{\mathrm{E},\mathrm{d}}/\eta^{\mathrm{d}})\Delta t=0,$ (17c)
$0\leq P^{\text{L,s}}_{t}\leq\rho\hat{P}^{\text{L}}_{t},$ (17d)
$P^{\text{EX,C}}_{t}=-(\hat{P}_{t}^{\text{R}}-P_{t}^{\text{R,c}})+\hat{P}^{\text{L}}_{t}-P^{\text{L,s}}_{t}+P^{\text{E,c}}_{t}-P^{\text{E,d}}_{t},$ (17e)
$\underline{P}^{\text{EX}}_{t}\leq P^{\text{EX,C}}_{t}\leq\overline{P}^{\text{EX}}_{t},$ (17f)
where (17c) denotes period constraint of the ES. $\alpha^{\text{L,s}}$ is the flexible load adjustment cost coefficient. $P^{\text{EX,C}}_{t}$ is the trading power of the commercial ECs.
-A3 Industrial ECs
Industrial ECs (e.g., manufacturing, mining, and construction industries) are typically equipped with PV, energy storage, inflexible load, flexible load, and distributed generator (DG). Considering the complex operational characteristics of industrial production processes, flexible loads exist in the form of discrete entire processes and can only be shifted as a whole in terms of temporal sequence.
The objective of industrial ECs operation is to minimize the operational cost including cost for power trading, flexible load adjustment, RES reduction and DG generation. Thus, the overall operation model of industrial ECs is as follows:
$\!\!\min$ $\sum_{t=1}^{T}\!\Delta t(\pi^{\text{EC}}_{t}P^{\text{EX,I}}_{t}\!+\!\alpha^{\text{DG}}P^{\text{DG}}_{t}\!+\!\sum_{i=1}^{N^{\mathrm{s}}}\!\alpha_{i}^{\text{L,s}}P^{\text{L,s}}_{i,t}\Delta t^{\text{L,s}}_{i}\!+\!\alpha^{\text{R,c}}P^{\text{R,c}}_{t}),$ (18a)
[…]
$0\leq P^{\text{DG}}_{t}\leq\overline{P}^{\text{DG}},$ (18c)
$-\Delta\overline{P}^{\text{DG}}\leq P^{\text{DG}}_{t}-P^{\text{DG}}_{t-1}\leq\Delta\overline{P}^{\text{DG}}_{t},$ (18d)
$\tilde{P}^{\text{L,s}}_{i,t-\Delta t_{i}^{\text{L,s}}}=P^{\text{L,s}}_{i,t},\ \forall t\in[t^{1}_{i},t^{2}_{i}],\ \forall i\in[1,N^{\mathrm{s}}],$ (18e)
$\tilde{P}^{\text{L}}_{t}=P^{\text{L}}_{t}-\sum_{i=1}^{N^{\mathrm{s}}}P^{\text{L,s}}_{i,t}+\sum_{i=1}^{N^{\mathrm{s}}}\tilde{P}^{\text{L,s}}_{i,t},$ (18f)
$P^{\text{EX,I}}_{t}=-(P_{t}^{\text{R}}-P_{t}^{\text{R,c}})+\tilde{P}^{\text{L}}_{t}+P^{\text{E,c}}_{t}-P^{\text{E,d}}_{t}-P^{\text{DG}}_{t},$ (18g)
$\underline{P}^{\text{EX}}_{t}\leq P^{\text{EX,I}}_{t}\leq\overline{P}^{\text{EX}}_{t},$ (18h)
where $P^{\text{DG}}_{t}$ is the DG output power, and $\overline{P}^{\text{DG}}/\Delta\overline{P}^{\text{DG}}$ denote its maximum output/ramping limit. $a,b$ are the DG cost coefficients. $\tilde{P}^{\text{L}}_{t}$ is the adjusted load demand. $\tilde{P}^{\text{L,s}}_{i,t}/\Delta t_{i}^{\text{L,s}}$ are the adjusted power and duration of the $i$-th flexible load ($N^{\mathrm{s}}$ in total), where $t^{1}_{i}/t^{2}_{i}$ denote the start/end time periods.

### v2

@@ V Conclusion
In this paper, a quantum learning and estimation approach is proposed to address the bilevel DNs-ECs coordinated operation problem. By leveraging quantum properties including quantum superposition and entanglement, the proposed Q-TCN-LSTM model can efficiently learn the incentive-response characteristics of ECs, realizing accurate estimation of user electricity consumption behavior under limited information. Additionally, QAE is introduced to facilitate evaluations with numerous scenarios, achieving a great speedup over classical MC methods, significantly reducing computational resource consumption and execution time. In this way, the proposed quantum learning and estimation approach efficiently facilitates the coordinated operation between DNs and ECs to generate better pricing strategies to regulate the net load profiles of ECs, thereby reducing the operational risk introduced to the DNs. As illustrated by the case studies, the proposed quantum-enhanced approach demonstrates superior performance in terms of both mapping accuracy and computational efficiency.
Although quantum learning and estimation methods have demonstrated certain theoretical advantages over their classical counterparts, the practical application of quantum computing in power systems is constrained by current quantum hardware limitations. First, hardware performance and scalability are inadequate. Existing quantum processors provide only a few hundred qubits with short coherence times and high noise sensitivity, limiting their ability to handle large-scale power system tasks. Although error correction may improve reliability, it requires significant additional resources. Second, the quantum–classical interface restricts efficiency. Classical data must be encoded into quantum states and later decoded through measurement, causing potential information loss and being limited by qubit availability. Third, quantum hardware is costly and difficult to integrate with existing infrastructures. Devices require extreme operating conditions and are incompatible with current classical data systems, such as SCADA, without substantial modifications. Hence, advancing practical quantum hardware is critical to converting theoretical advantages into engineering value, enabling accurate risk assessment and optimized operation in power systems.